OpenStudy (anonymous):

Solve x2 - 7x = -13.

OpenStudy (anonymous):

by completing the square and using the quadratic formula

OpenStudy (ranga):

x^2 - 7x = -13 (x - 7/2)^2 = -13 + (7/2)^2 Simplify the right. Take square roots on both sides and solve for x.

OpenStudy (ranga):

To complete the square for the general quadratic equation: ax^2 + bx + c = 0 Make the leading coefficient 1 by dividing throughout by a x^2 +(b/a)x + (c/a) = 0 Take the constant to the other side x^2 +(b/a)x = -(c/a) { x + (b/(2a)) }^2 = -(c/a) + (b/(2a))^2 So b/2a will be the term that will go inside the square.